# Improving planning and MBRL with temporally-extended actions

### v1

@@ Abstract
Continuous time systems are often modeled using discrete time dynamics but this requires a small simulation step to maintain accuracy. In turn, this requires a large planning horizon which leads to computationally demanding planning problems and reduced performance. Previous work in model free reinforcement learning has partially addressed this issue using action repeats where a policy is learned to determine a discrete action duration. Instead we propose to control the continuous decision timescale directly by using temporally-extended actions  and letting the planner treat the duration of the action as an additional optimization variable along with the standard action variables. This additional structure has multiple advantages. It speeds up simulation time of trajectories and, importantly, it allows for deep horizon search in terms of primitive actions while using a shallow search depth in the planner. In addition, in the model based reinforcement learning (MBRL) setting, it reduces compounding errors from model learning and improves training time for models. We show that this idea is effective and that the range for action durations can be automatically selected using a multi-armed bandit formulation and integrated into the MBRL framework. An extensive experimental evaluation both in planning and in MBRL, shows that our approach yields faster planning, better solutions, and that it enables solutions to problems that are not solved in the standard formulation.
@@ 1 Introduction
Shooting based planners like Cross Entropy Method (CEM) and Model Predictive Path Integral (MPPI) (Kobilarov, 2012; Botev et al., 2013; Williams et al., 2017; Chua et al., 2018) rely on sampling to estimate the highly rewarding regions of the state space. They are known to perform poorly as the planning horizon increases, especially in environments with noisy dynamics or sparse rewards. This is because in such environments the variance of their estimates increase and they need a large number of samples in order to act reliably (Chatterjee et al., 2023). In addition, when planning with learned models as in MBRL, longer rollouts can lead to compounding errors (Janner et al., 2019).
@@
In this paper, we propose to control the decision timescale directly by treating the action duration $(\delta t)$ as an additional optimization variable. At each decision point, the planner optimizes for the standard primitive action as well as its duration resulting in a temporally-extended action. In contrast to using a fixed frame-skip, the duration for each temporally-extended action  can be different allowing the planner to choose actions of varying durations, as can be seen in Figure 4. This provides the planner with an added degree of flexibility and, at the same time, it reduces the search space by constraining the structure of the trajectories that the planner searches over. Further, as a small increase in the temporally-extended  planning horizon can translate to a large increase in the primitive planning horizon, this allows the agent to search deeper, enabling it to solve environments which are otherwise too difficult.
Finally, unlike previous work, we do not learn a policy. Our work is in the context of MBRL. We learn a model that predicts the next state and reward due to a temporally-extended action  and use the learned model to plan, leading to superior performance as well as faster training. Selecting a suitable range for $\delta t$ is an important step as it can directly impact the performance of the planner. We show that posing this as a non-stationary multi-armed bandit (MAB) (Besbes et al., 2014; Lattimore and Szepesvári, 2020) problem allows us to select the range automatically.
@@ 2 Related Work
Frame-skips or action-repeats are discrete values and, while they introduce a temporal abstraction for discrete time systems, a continuous representation is both more realistic and more flexible. Some prior work in model-free RL has explored this problem. Ni and Jang (2022) use a modified Soft Actor Critic (Haarnoja et al., 2018) to learn a policy and control the timescale rather than the action duration, with an explicit constraint on the average timescale of the policy. Their formulation of the optimization objective is similar to ours except that they use a linear approximation of the reward function to compute the reward due to a macro-action (details in Appendix B). Wang and Beltrame (2024) introduce Soft-Elastic Actor Critic to output the duration of the action along with the action itself. However, they modify the reward structure by introducing penalties for the energy and the time taken by each action. Our work is in the planning and MBRL setting. Rather than learn a policy, we use either exact or learned transition and reward functions to plan. Our objective arises naturally from the formulation without the need to add additional constraints or engineer rewards. Further, in previous work in model-free RL, the maximum value of action-repeats or timescale is set as a hyperparameter. Instead, we use a MAB framework, similar to Li et al. (2018) and Lu et al. (2022), and automatically select the maximum action duration removing the need for tuning an additional hyperparameter.
@@ 3 Modeling Temporally-Extended Actions
Let us assume that the agent has access to the primitive dynamics function $(f)$ which is accurate for all $0\leq t\leq\delta_{t}^{\text{env}}$. In the standard setup, the agent uses $f$ at each decision time and outputs an action whose duration is implicitly $\delta_{t}^{\text{env}}$. The return due to a trajectory $\tau$ is given by $J_{1}=\sum_{t=1}^{L(\tau)}\gamma^{t-1}\mathcal{R}(s_{t},a_{t})$ where $L(\tau)$ is the number of decision points in $\tau$. Here, the number of execution steps is exactly equal to the number of decision steps.
@@
where $e_{<1}=0$. This formulation avoids the need for a precise continuous time model of discounting. Note that when $\delta t_{k}=\delta_{t}^{\text{env}}$ for all $k$, then $J_{2}=J_{1}$.
@@
Algorithm 1 One decision step : Action selection using a temporally-extended dynamics function ($F$) with a shooting-based planner
1:temporally-extended  dynamics function $(F)$, current state $(s_{k})$
@@
3:     Sample $N$ action sequences of length $D_{\text{TE}}$ using $\mu_{a},\text{var}_{a}$
4:     for all $N$ action sequences $(a_{t},\dots,a_{t+D_{\text{TE}}})$ do
5:         $//~{}a_{t}$ includes standard action variables along with their duration
6:         for step $t=0$ to $D_{\text{TE}}-1$ do
7:              Simulate transition using $F$ and $a_{t}$ and compute aggregate reward
8:         end for
9:     end for
10:     Update $\mu_{a},\text{var}_{a}$ using action sequence and aggregate reward
11:end for
12:return sample $a_{k}$ using $\mu_{a},\text{var}_{a}$
@@ 4 Method
We want the planner to have access to a temporally-extended dynamics function $(F)$ that can work with temporally-extended actions. If $F$ is available, then using it for planning is straightforward. For example, one can use a shooting-based planner with $F$ to select an action from state $s_{k}$ as shown in Algorithm 1. However, $F$ is usually not readily available. In the standard planning setup, the agent has access to the primitive dynamics function $(f)$ which samples from one step primitive transition and reward function. As a simple solution, we can wrap $f$ in a loop (as shown in Algorithm A1 in the Appendix) to obtain $F_{\text{IP}}$. This holds as we assumed $f$ is accurate for all $0\leq t\leq\delta_{t}^{\text{env}}$. We call this the iterative primitive dynamics function. Although $F_{\text{IP}}$ accurately captures $F$, it fails to facilitate the speed of execution. The time required by $F_{\text{IP}}$ to simulate the outcome due to a temporally-extended action  is dependent on the duration of the action itself. However, if we had access to $F$, this would have been a constant time evaluation.
To fix this, we use neural networks and approximate $F$ using $\hat{F}_{\text{TE}}$. We can then use $\hat{F}_{\text{TE}}$ to predict the next state and reward due to a temporally-extended action from a given state. For learning $\hat{F}_{\text{TE}}$, we use the framework of MBRL. We collect data by interacting with the environment and use the data to train a neural network to predict a distribution over the next states and a point estimate for the reward. Once $\hat{F}_{\text{TE}}$ is learned, it can be used for planning. In this work, we use a shooting-based planner and use Algorithm 1 with $\hat{F}_{\text{TE}}\approx F$ in order to select actions.
@@
We set $\delta t_{\min}=\delta_{t}^{\text{env}}$ to be the timescale of the environment. For $\delta t_{\max}$, we consider exponentially-spaced candidates, and pose the selection as a MAB problem, where each $\delta t_{\max}$ candidate is modeled as an arm. Pulling an arm is equivalent to selecting one of the $\delta t_{\max}$ candidates and using it to collect data for one episode. The aggregate reward observed at the end of an episode indicates the quality of the arm. Note that the rewards obtained are directly related to the quality of the learned dynamics model the agent has. As the agent collects more data, the model quality and hence the reward distribution of the arms change, making the problem of arm selection, non-stationary. To address this we propose to use an exponential moving average that focuses the reward estimate on recent episodes and hence ameliorates the effect of non-stationarity. Another important question is whether to learn a single dynamics model to be shared by all the arms or learn a separate dynamics model for each arm (see Section H.1 for a larger discussion). We find that learning a separate dynamics model for each arm and using a UCB heuristic with exponential moving average of the rewards per timestep works well for our purpose.
Let $\bar{R}_{T}$ be the average reward per timestep obtained in episode $T$ and let $\hat{R}_{i,T}=\hat{R}_{i,T-1}+\alpha(\bar{R}_{T}-\hat{R}_{i,T-1})$ be the exponentially-weighted average of rewards per timestep obtained till episode $T$ due to candidate $i$. Let $N(i,T)$ be the number of times candidate $i$ was chosen before iteration $T$. At the start of iteration $T+1$, $\delta t_{\max}$ is selected using Equation 3 where we set $\hat{R}_{i,T=0}$ to the mean reward obtained across 5 episodes using the randomly initialized dynamics models, prior to any training.
@@ 5 Experiments
We evaluate the use of temporally-extended actions both in planning and in MBRL. We compare the planning performance of $A_{\text{STD}}$ and $A_{\text{TE}}$ using CEM, which is a shooting-based planner, as used previously by Chua et al. (2018). CEM maintains a sequence of sampling distributions from which it generates multiple action sequences. For each action sequence, it instantiates multiple particles and computes the trajectory and reward due to each particle. Then, it computes the mean reward per action sequence and uses the top $k$ action sequences to bias its sampling distributions. For evaluation, we wrap the primitive dynamics function of the simulation environment in an iterative loop as shown in Algorithm A1.
@@
[…]
We consider two variants of our proposed algorithm depending on whether the range of $\delta t$, specified by $\delta t_{\min}$ and $\delta t_{\max}$, is fixed or selected dynamically at the start of each episode. $A_{\text{TE}}(\text{F})$ uses a fixed value of $\delta t_{\max}$, while $A_{\text{TE}}(\text{D})$ uses the proposed MAB framework to dynamically select $\delta t_{\max}$ at the beginning of each episode. Both variants use $\delta t_{\min}=\delta_{t}^{\text{env}}$. The other key difference is that $A_{\text{TE}}(\text{D})$ learns a temporally-extended dynamics model for each $\delta t_{\max}$ candidate. For our discussion, we use $A_{\text{TE}}$ to refer to both the variants in general while specifying the variant wherever required. Full details of the experimental setup are given in Appendix D.
First, we experiment in a planning regime where the agent has access to the exact dynamics function. In this setting, $A_{\text{TE}}$ uses iterative primitive dynamics function (Algorithm A1) for planning. For experiments, we use the Mountain Car environment from Gymnasium (Kwiatkowski et al., 2024), a multi-hill Mountain Car environment from the Probabilistic and Reinforcement Learning Track of the International Planning Competition (IPC) 2023 (Taitler et al., 2024), and the Dubins car environment from Chatterjee et al. (2023). Then, we experiment in the MBRL setting where $\mathcal{T}$ and $\mathcal{R}$ are not known. In this case, we learn a temporally-extended model as discussed above, by interacting with the environment. We use Cartpole from Gymnasium and Ant, Half Cheetah, Hopper, Reacher, Pusher and Walker from MuJoCo (Todorov et al., 2012). We run each experiment across 5 different seeds and aggregate the results.
Does planning with temporally-extended actions help? Mountain Car has a sparse reward, requiring a large planning horizon to succeed. We compare the performance of $A_{\text{STD}}$ and $A_{\text{TE}}$ by varying the planning horizon. As shown in Figure 1, $A_{\text{TE}}(D_{\text{TE}}\geq 4)$ solves the environment with a much smaller planning horizon than $A_{\text{STD}}(D_{\text{STD}}\geq 60)$.
Next, we experiment with the multi-hill version of Mountain Car from IPC 2023 (illustrated in Figure 1(b)). Each instance of the environment increases the difficulty by either adding more hills or altering the surface of the hills. As shown in Table 1, while $A_{\text{STD}}$ is able to solve only the first instance, $A_{\text{TE}}$ solves all the instances. Note that, even though $A_{\text{TE}}$ takes a small number of decision steps, the computation time required to finish an episode is comparable to $A_{\text{STD}}$. This is because $A_{\text{TE}}$ uses an iterative version of the dynamics $(F_{\text{IP}})$ for simulation. Another observation is that as the episode terminates when the agent reaches the goal, the number of decision steps is smaller than $D_{\text{TE}}$.
@@
Can temporally-extended actions help transform infeasible problems to feasible ones? We use the Dubins Car environment $(\delta_{t}^{\text{env}}=0.2)$ and experiment with u-shaped map as shown in Figure 2. This is a challenging configuration, where shooting-based planners often fail, because the car is initially facing the obstacles and a naive forward search hits the obstacles and does not yield useful information. In addition, as the reward is sparse, this map requires the planning horizon to be large. To solve the environment, $A_{\text{STD}}$ requires 10,000 samples with a planning horizon of 1000, while $A_{\text{TE}}$ requires a planning horizon of 75 and $\delta t_{\max}=20$ (see detailed discussion in Section F.2).
To increase difficulty, we augment the action space with 100 dummy action variables. Although these variables do not contribute to the dynamics or rewards, the agent is unaware of this and has to account for all the action variables. Even in this setting, $A_{\text{TE}}$ succeeds while $A_{\text{STD}}$ fails due to large memory requirements. A simple computation shows that $A_{\text{STD}}$ needs around 4GB of memory to track the sampled actions, whereas $A_{\text{TE}}$ requires just 103MB. In other configurations that reduce the memory requirements (detailed in Section F.2), the search fails to find the goal. This shows that using temporally-extended actions can often turn infeasible search problems into feasible ones.
@@
How does varying the two discount factors impact the agent’s behaviour? The experiments so far use a single value of $\gamma$ while the proposed formulation in Equation 2 has two discount factors. As we discuss next, while some expectations on the effect of $\gamma_{1}$ and $\gamma_{2}$ are intuitive, a complete characterization is not obvious.
@@
To explore the impact of $\gamma_{1}$ and $\gamma_{2}$, we fix one and vary the other. We use the cave-mini map (Figure 1(a)) in the Dubins Car environment, where the agent has to navigate through multiple obstacles. We have two hypotheses. First, for a fixed $\gamma_{1}$, smaller values of $\gamma_{2}$ will prefer shorter action durations and hence larger number of decision points. Second, for a fixed $\gamma_{2}$, decreasing $\gamma_{1}$ should prefer longer action durations and a smaller total number of primitive actions. The intuition is that if we reduce $\gamma_{1}$ and the planner does not reduce the number of primitive actions, then the impact of discounting on the overall objective will be higher. Table 2 confirms both these trends.
@@
Each experiment is run for a specific number of iterations, where in each iteration, we train the model using multiple mini-batches drawn from the dataset of past transitions, and then use the trained model to act in the environment for one episode. The rewards collected during the episode are used for evaluating the agent’s performance. We plot the mean and standard deviation of the running average of scores obtained as well as the number of decisions made by the agents across training iterations. The main results from our experiments are in Figure 3.
Overall, $A_{\text{TE}}$ performs better than $A_{\text{STD}}$ while being faster. $A_{\text{TE}}(\text{F})$, with a suitable selection of $\delta t_{\max}$, results in the best performance. $A_{\text{TE}}(\text{D})$ requires more training iterations than $A_{\text{TE}}(\text{F})$ as it has to spend time exploring all the candidate arms but the performance catches up. In addition, $A_{\text{TE}}$ is significantly faster than $A_{\text{STD}}$ in terms of computation time (see Table A3 in the Appendix). The speed-up for $A_{\text{TE}}$ comes from three sources. First, using temporally-extended actions  leads to fewer decision points. Second, the reduction in search space, due to temporally-extended actions  further reduces planning time. Third, a decrease in decision points results in a smaller sized dataset of past transitions that the model needs to learn from, thus reducing the model training time.

### v2

@@ Abstract
Continuous time systems are often modeled using discrete time dynamics but this requires a small simulation step to maintain accuracy. In turn, this requires a large planning horizon which leads to computationally demanding planning problems and reduced performance. Previous work in model-free reinforcement learning has partially addressed this issue using action repeats where a policy is learned to determine a discrete action duration. Instead we propose to control the continuous decision timescale directly by using temporally-extended actions  and letting the planner treat the duration of the action as an additional optimization variable along with the standard action variables. This additional structure has multiple advantages. It speeds up simulation time of trajectories and, importantly, it allows for deep horizon search in terms of primitive actions while using a shallow search depth in the planner. In addition, in the model-based reinforcement learning (MBRL) setting, it reduces compounding errors from model learning and improves training time for models. We show that this idea is effective and that the range for action durations can be automatically selected using a multi-armed bandit formulation and integrated into the MBRL framework. An extensive experimental evaluation both in planning and in MBRL, shows that our approach yields faster planning, better solutions, and that it enables solutions to problems that are not solved in the standard formulation.
@@ 1 Introduction
Shooting based planners like Cross Entropy Method (CEM) and Model Predictive Path Integral (MPPI) (Kobilarov, 2012; Botev et al., 2013; Williams et al., 2017; Chua et al., 2018) rely on sampling to estimate the highly rewarding regions of the state space. They are known to perform poorly as the planning horizon increases, especially in environments with noisy dynamics or sparse rewards. This is because in such environments the variance of their estimates increases and they need a large number of samples in order to act reliably (Chatterjee et al., 2023). In addition, when planning with learned models as in MBRL, longer rollouts can lead to compounding errors (Janner et al., 2019).
@@
In this paper, we propose to control the decision timescale directly by treating the action duration $(\delta t)$ as an additional optimization variable. At each decision step, the planner optimizes for the standard primitive action as well as its duration resulting in a temporally-extended action. In contrast to using a fixed frame-skip, the duration for each temporally-extended action  can be different allowing the planner to choose actions of varying durations, as can be seen in Figure˜4. This provides the planner with an added degree of flexibility and, at the same time, it reduces the search space by constraining the structure of the trajectories that the planner searches over. Further, as a small increase in the temporally-extended  planning horizon can translate to a large increase in the primitive planning horizon, this allows the agent to search deeper, enabling it to solve environments which are otherwise too difficult.
Finally, unlike previous work, we do not learn a policy. Our work is in the context of MBRL. We learn a model that predicts the next state and reward due to a temporally-extended action  and use the learned model to plan, leading to superior performance as well as faster training. Selecting a suitable range for $\delta t$ is an important step as it can directly impact the performance of the planner. We show that posing this as a non-stationary multi-armed bandit (MAB) problem (Besbes et al., 2014; Lattimore and Szepesvári, 2020) allows us to select the range automatically.
@@ 2 Related Work
Frame-skips or action-repeats are discrete values and, while they introduce a temporal abstraction for discrete time systems, a continuous representation is both more realistic and more flexible. Some prior work in model-free RL has explored this problem. Ni and Jang (2022) use a modified Soft Actor Critic (Haarnoja et al., 2018) to learn a policy and control the timescale rather than the action duration, with an explicit constraint on the average timescale of the policy. Their formulation of the optimization objective is similar to ours except that they use a linear approximation of the reward function to compute the reward due to a macro-action (details in Appendix˜B). Wang and Beltrame (2024) introduce Soft-Elastic Actor Critic to output the duration of the action along with the action itself. However, they modify the reward structure by introducing penalties for the energy and the time taken by each action. Our work is in the planning and MBRL setting. Rather than learn a policy, we use either exact or learned transition and reward functions to plan. Our objective arises naturally from the formulation without the need to add additional constraints or engineer rewards. Further, previous work in model-free RL treat the maximum value of action-repeats or timescale as a hyperparameter. Instead, we use a MAB framework, similar to Li et al. (2018) and Lu et al. (2022), and automatically select the maximum action duration removing the need for tuning an additional hyperparameter.
@@ 3 Modeling Temporally-Extended Actions
Let us assume that the agent has access to the primitive dynamics function $(f)$ which is accurate for all $0\leq t\leq\delta_{t}^{\text{env}}$. In the standard setup, the agent uses $f$ at each decision step and outputs an action whose duration is implicitly $\delta_{t}^{\text{env}}$. The return due to a trajectory $\tau$ is given by $J_{1}=\sum_{t=1}^{L(\tau)}\gamma^{t-1}\mathcal{R}(s_{t},a_{t})$ Here, the number of execution steps is exactly equal to the number of decision steps.
@@
where $e_{<1}=0$. This formulation provides an approximation that avoids the need for a precise continuous time model of discounting. Note that when $\delta t_{k}=\delta_{t}^{\text{env}}$ for all $k$, then $J_{2}=J_{1}$.
@@
Algorithm 1 One decision step : Action selection using a temporally-extended dynamics function ($F$) with a generic shooting-based planner
1:temporally-extended  dynamics function $(F)$, current state $(s)$
@@
3:  Sample NN action sequences of length $D_{\text{TE}}$ using $\mu_{a},\text{var}_{a}$
4:  for all NN action sequences $\textbf{a}_{t:t+D_{\text{TE}}}=(a_{t},\dots,a_{t+D_{\text{TE}}})$ do
5:   $a_{t}$ includes standard action variables and the action duration
6:   Simulate trajectory using $F$ from $s$ due to $\textbf{a}_{t:t+D_{\text{TE}}}$ and compute aggregate reward
7:  end for
8:  //update rule is specific to the planner being used.
9:  Update $\mu_{a},\text{var}_{a}$ using action sequences and aggregate rewards
10:end for
11:return sample action using $\mu_{a},\text{var}_{a}$
@@ 4 Method
We want the planner to have access to a temporally-extended dynamics function $(F)$ that can work with temporally-extended actions. If $F$ is available, then using it for planning is straightforward. For example, one can use a shooting-based planner with $F$ to select an action from any state $s$ as shown in Algorithm˜1. However, $F$ is usually not readily available. In the standard planning setup, the agent has access to the primitive dynamics function $(f)$ which samples from the one step primitive transition and reward distribution. A simple solution is to wrap $f$ in a loop (as shown in Algorithm˜A1 in the Appendix) to obtain $F_{\text{IP}}$. This holds as we assumed that $f$ is accurate for all $0\leq t\leq\delta_{t}^{\text{env}}$. We call this the iterative primitive dynamics function. Although $F_{\text{IP}}$ accurately captures $F$, it fails to facilitate the speed of execution. The time required by $F_{\text{IP}}$ to simulate the outcome due to a temporally-extended action  is dependent on the duration of the action itself. However, if we had access to $F$, this would have been a constant time evaluation.
Approximating $F$ : We address this by using neural networks to approximate $F$ using $\hat{F}_{\text{TE}}$. This allows us to predict the next state and reward due to a temporally-extended action from a given state in constant time. For learning this model, we use the framework of MBRL. We collect data by interacting with the environment and use the data to train $\hat{F}_{\text{TE}}$ to predict a distribution over the next states and a point estimate for the reward. The learned model can then be used for planning. For example, Algorithm˜1 can be used with $\hat{F}_{\text{TE}}$ instead of $F$ to select actions, as we do in this work.
@@
We set $\delta t_{\min}=\delta_{t}^{\text{env}}$ to be the timescale of the environment. For $\delta t_{\max}$, we consider $m$ exponentially-spaced candidates, where $m=\log_{2}(T)$ and $T$ is the maximum number of primitive steps defined by the environment. We model each $\delta t_{\max}$ candidate as an arm and pose the selection as a MAB problem. Pulling an arm is equivalent to selecting one of the $\delta t_{\max}$ candidates and using the planner with this value of $\delta t_{\max}$ to collect data for one episode. The aggregate reward observed at the end of an episode indicates the quality of the arm. However, note that the rewards obtained are directly related to the quality of the learned dynamics model the agent has. As the agent collects more data, the model quality and hence the reward distribution of the arms change, making the problem of arm selection non-stationary. To address this we propose to use an exponential moving average that focuses the reward estimate on recent episodes and hence ameliorates the effect of non-stationarity.
Let $\bar{R}_{T}$ be the average reward per timestep obtained in episode $T$ and let $\hat{R}_{i,T}=\hat{R}_{i,T-1}+\alpha(\bar{R}_{T}-\hat{R}_{i,T-1})$ be the exponentially moving average (EMA) of rewards per timestep obtained till episode $T$ due to candidate $i$. Let $N(i,T)$ be the number of times candidate $i$ was chosen before iteration $T$. At the start of iteration $T+1$, $\delta t_{\max}$ is selected using the UCB heuristic (Auer et al., 2002) in Equation˜3 where we set $\hat{R}_{i,T=0}$ to the mean reward obtained across 5 episodes using the randomly initialized dynamics models, prior to any training.
@@
Another important question is whether to learn a single dynamics model to be shared by all the arms or learn a separate dynamics model for each arm. Our preliminary experiments, which we describe in Section˜I.1, suggested that learning a separate dynamics model for each $\delta t_{\max}$ candidate is better.
To summarize, we treat each $\delta t_{\max}$ candidate as a bandit arm. Before each episode, we use a UCB heuristic with EMA of rewards to select an arm and use the selected value of $\delta t_{\max}$ and the corresponding dynamics model to plan for the duration of the episode.
@@ 5 Experiments
We evaluate the use of temporally-extended actions in planning and in MBRL. We begin by elaborating on our method for both of these settings, before discussing the experiments and results.
Experimental setup : We compare the planning performance of $A_{\text{STD}}$ and $A_{\text{TE}}$ using CEM, which is a shooting-based planner. CEM maintains a sequence of sampling distributions from which it generates multiple action sequences. For each action sequence, it instantiates multiple particles and computes the trajectory and reward due to each particle. Then, it computes the mean reward per action sequence and uses the top $k$ action sequences to bias its sampling distributions. For evaluation, we wrap the primitive dynamics function of the simulation environment in an iterative loop using $F_{\text{IP}}$ as discussed in the previous section.
[…]
Algorithm 2 Using $A_{\text{TE}}$ with a fixed $\delta t_{\max}$ as planner for MBRL
[…]
2:Initialize empty dataset $\mathcal{D}$
3:for $i=1$ to $n$ do
4:  //data collection
5:  for $t=1$ to end of episode do
6:   $a_{t}$ includes standard action variables and the action duration
7:   $a_{t}$ = planner.choose_action($s_{t}$, $\delta t_{\max}$)
8:   //env.step() implicitly uses temporally-extended dynamics
9:   $s_{t+1},r_{t}$ = env.step($a_{t}$)
10:   $\mathcal{D}$ = $\mathcal{D}$.append(($s_{t},a_{t},s_{t+1},r_{t}$))
11:  end for
12:  //training
13:  num_of_batches = len($\mathcal{D}$) / $b$
14:  for $j=1$ to num_of_batches do
15:   for each dynamics model in $\hat{F}_{\text{TE}}$ do
16:     Sample a mini-batch of size $b$ from $\mathcal{D}$
17:     Compute the loss and gradients using the mini-batch
18:     Update the model using the gradients
19:   end for
20:  end for
21:end for
We consider two variants of our proposed algorithm - $A_{\text{TE}}(\text{F})$ which uses a fixed and manually chosen value of $\delta t_{\max}$ and $A_{\text{TE}}(\text{D})$ which uses the proposed MAB framework to dynamically select the value of $\delta t_{\max}$ at the beginning of each episode. Both variants use $\delta t_{\min}=\delta_{t}^{\text{env}}$. A key difference between the two is that $A_{\text{TE}}(\text{D})$ maintains a separate dataset and learns a separate temporally-extended dynamics model for each $\delta t_{\max}$ candidate. Specifically, $A_{\text{TE}}(\text{F})$ uses Algorithm˜2 exactly, while $A_{\text{TE}}(\text{D})$ has an additional step where it uses the MAB framework to select a $\delta t_{\max}$ value before the start of each episode and then uses the corresponding dataset and model for the duration of the episode. For our discussion, we use $A_{\text{TE}}$ to refer to both the variants in general while specifying the variant wherever required.
Experimental Evaluation
First, we experiment in a planning regime where the agent has access to the exact dynamics function. In this setting, $A_{\text{TE}}$ uses iterative primitive dynamics function (Algorithm˜A1) for planning. For experiments, we use the Mountain Car environment from Gymnasium (Kwiatkowski et al., 2024), a multi-hill Mountain Car environment from the Probabilistic and Reinforcement Learning Track of the International Planning Competition (IPC) 2023 (Taitler et al., 2024), and the Dubins car environment from Chatterjee et al. (2023). Then, we experiment in the MBRL setting where $\mathcal{T}$ and $\mathcal{R}$ are not known. In this case, we learn a temporally-extended model as discussed above, by interacting with the environment. We use Cartpole from Gymnasium and Ant, Half Cheetah, Hopper, Reacher, Pusher and Walker from MuJoCo (Todorov et al., 2012). We run each experiment across 5 different seeds and aggregate the results. Full details of the experimental setup are given in Appendix˜E.
Does planning with temporally-extended actions help? Mountain Car has a sparse reward, requiring a large planning horizon to succeed. We compare the performance of $A_{\text{STD}}$ and $A_{\text{TE}}$ by varying the planning horizon. As shown in Figure˜1, $A_{\text{TE}}(D_{\text{TE}}\geq 4)$ solves the environment with a much smaller planning horizon than $A_{\text{STD}}(D_{\text{STD}}\geq 60)$.
@@
Next, we experiment with the multi-hill version of Mountain Car from IPC 2023 (illustrated in Figure˜1(b) in the Appendix). Each instance of the environment increases the difficulty by either adding more hills or altering the surface of the hills. As shown in Table˜1, while $A_{\text{STD}}$ is able to solve only the first instance, $A_{\text{TE}}$ solves all the instances. Note that, even though $A_{\text{TE}}$ takes a small number of decision steps, the computation time required to finish an episode is comparable to $A_{\text{STD}}$. This is because $A_{\text{TE}}$ uses an iterative version of the dynamics $(F_{\text{IP}})$ for simulation. Another observation is that because the episode terminates when the agent reaches the goal, the number of decision steps is smaller than $D_{\text{TE}}$.
@@
Can temporally-extended actions help transform infeasible problems to feasible ones? We use the Dubins Car environment $(\delta_{t}^{\text{env}}=0.2)$ and experiment with u-shaped map as shown in Figure˜2. This is a challenging configuration, where shooting-based planners often fail, because the car is initially facing the obstacles and a naive forward search hits the obstacles and does not yield useful information. In addition, as the reward is sparse, this map requires the planning horizon to be large. To solve the environment, $A_{\text{STD}}$ requires 10,000 samples with a planning horizon of 1000, while $A_{\text{TE}}$ requires a planning horizon of 75 and $\delta t_{\max}=20$ (see detailed discussion in Section˜G.2).
To increase difficulty, we augment the action space with 100 dummy action variables. Although these variables do not contribute to the dynamics or rewards, the agent is unaware of this and has to account for all the action variables. We show that in this difficult setting as well, $A_{\text{TE}}$ succeeds while $A_{\text{STD}}$ fails due to large memory requirements. A simple computation shows that $A_{\text{STD}}$ needs around 4GB of memory to track the sampled actions, whereas $A_{\text{TE}}$ requires just 103MB. In other configurations that reduce the memory requirements (detailed in Section˜G.2), the search of $A_{\text{STD}}$ fails to find the goal. This shows that using temporally-extended actions can often turn infeasible search problems into feasible ones.
@@
How does varying the two discount factors impact the agent’s behaviour? The experiments so far use a single value of $\gamma$ while the proposed formulation in Equation˜2 has two discount factors. As we discuss next, while some expectations on the effect of $\gamma_{1}$ and $\gamma_{2}$ are intuitive, a complete characterization is not obvious.
@@
To explore the impact of $\gamma_{1}$ and $\gamma_{2}$, we fix one and vary the other. We use the cave-mini map (Figure˜1(a)) in the Dubins Car environment, where the agent has to navigate through multiple obstacles. We have two hypotheses. First, for a fixed $\gamma_{1}$, smaller values of $\gamma_{2}$ will prefer shorter action durations and hence larger number of decision steps. Second, for a fixed $\gamma_{2}$, decreasing $\gamma_{1}$ should prefer longer action durations and a smaller total number of primitive actions. The intuition is that if we reduce $\gamma_{1}$ and the planner does not reduce the number of primitive actions, then the impact of discounting on the overall objective will be higher. Table˜2 confirms both these trends.
@@
Each experiment is run for a specific number of iterations, where in each iteration, we train the model using multiple mini-batches drawn from the dataset of past transitions, and then use the trained model to act in the environment for one episode. The rewards collected during the episode are used for evaluating the agent’s performance. We plot the mean and standard deviation, across 5 different seeds, of the running average of scores obtained as well as the number of decisions made by the agents across training iterations. The main results from our experiments are shown in Figure˜3.
First, we review the results across environments before discussing environment specific observations. Overall, $A_{\text{TE}}$ performs better than $A_{\text{STD}}$ while being faster. $A_{\text{TE}}(\text{F})$, with a suitable selection of $\delta t_{\max}$, results in the best performance. $A_{\text{TE}}(\text{D})$ requires more training iterations than $A_{\text{TE}}(\text{F})$ as it has to spend time exploring all the candidate arms but it is able to catch up to the performance of $A_{\text{TE}}(\text{F})$ which uses a carefully selected value of $\delta t_{\max}$. In addition, $A_{\text{TE}}$ is significantly faster than $A_{\text{STD}}$ in terms of computation time (see Table˜A3 in the Appendix). The speed-up for $A_{\text{TE}}$ comes from three sources - (i) using temporally-extended actions leads to fewer decision points, (ii) the reduction in search space, due to temporally-extended actions further reduces planning time and, (iii) a decrease in decision points results in a smaller sized dataset of past transitions that the model needs to learn from, thus reducing the model training time.